# Assessing AI vs Human-Authored Spear Phishing SMS Attacks: An Empirical Study Using the TRAPD Method

### v1

@@ 5.2.7 Plausible Rewards
Realistic rewards significantly influenced the perceived credibility of phishing messages, as noted by 28% of the participants (7 out of 25). One participant said they would be more likely to click on a link, “because the reward is connected to something I put so much time and effort into” (T31), while another participant stressed the persuasiveness of plausible or realistic offers, mentioning, “It’s like taking you to a link to look at an offer, but it’s not a crazy insane offer” (T16). Some targets expressed enthusiasm for realistic opportunities, stating, “It’s something I’ll be really excited about, but also something that I think is realistic” (T34). Conversely, 24% of participants (6 out of 25) identified messages with rewards that seemed “too good to be true” as less convincing. One participant mentioned it’s “probably not likely” to be the “developer of the month four times a year” (T37). One of the targets questioned the credibility of messages offering free access to goods and services, stating, “It sounds too good to be true” (T34). Another target warned against messages promoting anything for free, noting, “Anything that’s free is a little too good to be true. Okay, so that’s when I would be very careful” (T27). This collective sentiment emphasized the convincing nature of realistic rewards in phishing messages while avoiding extravagant or implausible claims that trigger suspicion.
@@ 5.2.8 Sender Familiarity
Participants described how the perceived credibility of phishing messages is tied to the sender’s identity, as noted by 7 out of 25 targets (28%). One participant said the message seemed more convincing because the sender “introduced herself as [their] neighbor” (T30). Another explained the importance of a personal introduction, stating, “The sender started with ’My name is John’” (T41), which helped the target to trust the message. Several participants expressed the importance of connecting with the senders of the messages. One target was convinced by the “actual company branding” that mimicked promotional texts they had previously received (T18). Another explained, “This could be someone from my ward [church congregation], wanting me to like check out some plants or something” (T31). Participants explained that when the message sender appears to be a legitimate, familiar source, messages seemed more credible.
Shifting towards participants’ suspicions, a substantial 68% (17 out of 25) highlighted the pivotal role of the sender in raising skepticism about received messages. One participant explained that messages appear suspicious when the introduction is not consistent with other messages from that same sender, saying, “James is my boss… it’d be weird that he was introducing himself that way… that doesn’t sound like my boss” (T31). Several participants described the need to understand how the sender unexpectedly obtained their phone number and the purpose for which the sender sent a message. One target noted, “So like, how did you get my number in the first place?” (T27), while another questioned, “Like why would Lowe’s have my phone number?” (T20). Participants described how recognition of the sender’s phone number is also an important component of message believability. One participant explained, “If it’s from a number I don’t recognize at all. Then yeah, probably the first thing I would ask is who it is. And then however, if the person is like, if their name is like the same as one of my friends, then I probably would open it” (T34). Participants described that messages lacking personal connections from the sender raised suspicions, with one stating, “I don’t feel like it’s anyone who’s personally connected with me; it just seems like spam, you know?” (T31). Being familiar with the message sender was described as a strong indicator in whether subjects believed the messages to be legitimate. It is important to note that none of the study text messages sent to participants indicated a source phone number. Rather, our focus was on the content of the messages.
@@ 5.3 Human/AI Source Identification
This subsection addresses RQ3, which asks how effective targets are at identifying AI-generated spear phishing SMS messages. The statistical analysis presented earlier showed that targets could not effectively identify AI-generated messages. The qualitative results presented here reinforce the difficulty of this task for targets. Although the targets described some of their reasoning behind their guesses on which messages are AI-generated (see following section), 12 out of 25 of the participants stated in some way that they were uncertain about their decisions, often relying on intuition rather than any specific criteria. One target remarked, “sometimes it’s a gut feeling maybe more than like a specific thing you’re looking for?” (T07), while another said it’s“just a feeling” (T02). Others attributed their lack of criteria to advancements in AI, saying that they,“don’t really, really know my criteria, because I know that AI is getting so good” (T04). Several targets openly admitted their lack of expertise, saying they “have no idea” (T15), and “actually I don’t know, I don’t know” (T23). Some were unsure if AI could personalize content or include icons in text (T02), while others conveyed concerns of AI’s competency, hoping that “the ones that are worse would hopefully be the AI ones” (T34). Overall, these responses show the difficulty the targets faced in distinguishing between AI and human-generated content. In general, they did not seem to have mental models that provided meaningful guidance in identifying AI-generated messages.
@@ 5.4 AI Identification Criteria
This subsection addresses RQ4 by identifying the criteria people use to try and identify AI-generated messages.
@@ 5.4.2 URL
5.4.3 Word Choice
The message word choice was mentioned by 24% (6/25) of the targets. Four of the targets claimed that the word choice sounded like AI, particularly the use of buzzwords and marketing words. One target explained that messages sound AI-generated when they use buzzwords that are associated with being “attention grabbers” because “AI would utilize those a lot” (T16). Another target described a message as “too specific” when it included words that humans “wouldn’t necessarily use” when describing instructional design (T18). A few of the targets agreed that some of the messages had sentences that a human would have worded differently (T09). Conversely, two targets gave evidence of human-sounding messages and both mentioned slang words. One target described “There’s so many slang terms on it, that it seems really human to me” (T34), while another defended the messages they thought to be human-created and explained, “oftentimes, it was just because I felt like the language seemed a little more slang-like” (T06). Messages with specific, marketing-like words were perceived as AI-written while messages with casual, slang words that sounded more conversational were perceived as human-written.
5.4.4 Style
The message style was mentioned by 40% (10/25) of the targets. Many targets noted that AI-generated messages tend to be, “pretty formal” (T21) or “overly informal or overly formal” (T09) while others said messages often appear “extremely generic sounding,” as if they were created from a prompt with specific parameters (T18). Additionally, the presence of “glaring flaws” such as typos and awkward phrasings were attributed to being human-made mistakes. Such common mistakes contrasted with the “more robotic” nature of AI messages (T08). Some targets observed that AI-generated messages might seem “too perfect,” such as refraining from casual texting language that humans often use, such as “RN” for “right now” (T34). One participant noted that the excessive use of exclamation points is what made them think the message is AI generated: “I mean, honestly, what’s driving me insane about all of these is exclamation points. Yeah. I’m like, why are there so many exclamation points all over?” (T07). In contrast, messages that gave a sense of urgency were more likely to be perceived as human-written, as one target explained: “Then I didn’t put this one because they were actually pressuring me to do it immediately. So I don’t think AI can do it” (T02). Overall, targets indicated that a message’s style played a role in their AI vs human generated message decisions, with AI messages being attributed to using a style that is overly formal, generic, or laden with exclamation points; and with human messages appearing more casual, urgent, and imperfect.
5.4.5 Length of Message
The length of message was mentioned by 16% (4/25) of the targets. Participants assumed messages made by AI would be longer than human made messages. One mentioned that they would have ““AI do the longer ones and I would write the shorter ones myself,” (T21) assuming those who created these messages would do the same. These four targets were correct in their hunch that AI messages would be longer. Across the messages, AI-generated messages have on average 41% higher character count per message, with 337.8 characters for AI, and 237.9 characters for human-authored messages.
5.4.6 Content/Structure
The structure of the message content was mentioned by 24% (6/25) of the targets. Participants pointed out various aspects of structure that hinted at AI involvement, including repetitive phrases and/or awkward sentence construction. For instance, one participant thought a message was AI generated because ““it looks like there’s a template. So there’s like a flow that you know, there’s a pattern you see it as maybe started by a human” (T41). Another noted instances where messages seemed generic or mass-produced, akin to “the email that is sent to everyone in the school” (T41). One also pointed out inconsistencies within messages, such as repetitive phrases or ““two different things, in the same message” (T04). One target also noted that AI-generated messages tended to be more wordy, containing “lots of filler phrases” compared to human-authored messages (T37).
5.4.7 Grammar/Spelling
5.4.8 Personalization
The message personalization was mentioned by 32% (8/25) of the targets. Five participants mentioned that a lack of personalization made them suspect the message was AI-generated. For instance, one target noted, “AI were the ones that were a little bit less personal” (T34), while another expressed doubt, saying, “There is no connection with me as a reader” (T41). Some targets believed the absence of personal pronouns or a personal introduction with the recipient’s name indicated AI authorship. One target observed, “All of these messages, but two, include my name” (T37), leading them to suspect those two were AI-generated. Conversely, four participants cited personalized messages as evidence of human authorship, providing similar reasons as those who suspected AI. This group of participants felt messages tailored to the recipient sounded human, with one target commenting, “it just sounds very personal” (T21). Additionally, recipients noted that human-sounding messages often began with an introduction of the sender, such as “this is Sarah from the BYU student instructor program” (T41). Overall, personalized messages were deemed more human-like, while those lacking personalization were associated with AI.
@@ 6 Discussion
A growing body of literature has found evidence that AI-generated content from recent LLMs can outperform human-created content in various domains [8, 19, 20, 21]. For example, Zhang and Gosline compared content generated by GPT-4 with similar content created by professional content creators in the advertising field, finding that “Content generated by generative AI and augmented AI is perceived as of higher quality than that produced by human experts and augmented human experts” [19]. Nisbett and Spaiser explored the use of GPT-3 in creating moral statements supporting climate action, concluding that “GPT-3-generated statements are generally more convincing than human-generated statements” [21]. Similarly, in the domain of political speech, Palmer and Spirling found that “LLMs are capable of producing human-style arguments for different positions on subjects as varied as abortion, guns, immigration, and organ donation” and they could “out-perform human authors, though it varies by topic” [20].
@@
Fewer studies have examined the ability of LLMs to create content tailored to individuals, such as we examine in this paper. Heiding et al. examined phishing messages created by humans (using the V-triad approach), GPT-4, both combined, and a control message from an existing phishing dataset [22]. They found that the human V-triad and the combined human V-triad and GPT-4 approaches led to the highest click-through rate in a field experiment, followed by GPT-4 and then the control group. [22]. Although they did not personalize these messages to each individual, they did personalize them to a specific university context. They also provide a cost analysis that demonstrates how cheap the creation of spear phishing messages can be. Hazell also points out that LLM-generated spear phishing messages can be “realistic” and “cost effective,” but does not provide systematic evidence comparing humans to AI, such as presented in this paper [8].
@@
Our study contributes to this body of literature by comparing the effectiveness of AI-created and human-created spear phishing messages tailored to an individual target. We found that AI generally outperformed humans in creating spear phishing messages; however, the difference was not statistically significant, in part due to the relatively low sample size. As presented in Section 5.1.1, there is an 80% probability that GPT-4 is at least as good at or better than humans at created these highly personalized spear phishing messages. While further research is justified in this area, we believe our findings are a clear indication of the capabilities of LLMs in generating plausible spear phishing messages. It is important to note that these results are based on a very simple prompt, which could likely be improved. Furthermore, LLMs continue to improve in their quality, suggesting that AI will likely perform better at this task in the future. However, despite the fact that our human spear phishing authors had been trained on techniques to create spear phishing, it is possible that they could have been trained better. For example, they were not trained using the V-triad approach used so successfully by Heiding et al. [22].
@@ 6.2 Characteristics of Convincing Spear Phishing Messages
Our statistical analysis found that targets considered job-related spear phishing messages twice as persuasive than messages related to hobbies or social media posts (38% intended click rate compared to 19% hobbie and 17% social media). This was true for both human and AI-generated messages. This suggests the need to be particularly vigilant for work-related spear phishing attacks.
@@
The qualitative findings from this study point to several phishing message features that make them more convincing than others. [9, 12] Our findings aligned with prior research, revealing that messages matching the receiver’s expectations in terms of the sender, context, and relevance of a message were more convincing; those with poor grammar and spelling were less believable; and messages relating to the recipient’s life were more persuasive.
@@
Though we found general alignment across participants in terms of the features that matter, we also found that in some cases, the features that cause one person to “avoid phishing emails makes another person fall for them,” findings that were recently reported [22]. This suggests that some features may work better for certain individuals than others - a level of personalization that we have not yet examined, but seems possible to implement in future AI-generated spear phishing attacks that have more information about targets’ preferences.
@@
A key focus of this paper was on personalization of messages, since these were spear phishing messages tailored to individuals. Personalizing messages can lead to more persuasive messages [23, 10, 6], but can also raise red flags when the personalization is even slightly off. In this study, over two thirds of participants were likely to believe messages that related to topics of personal interest and relevance, showing the power of personalized phishing messages when done well. However, two thirds of participants also shared examples of messages that were not believed because they included content that was not personally relevant or included red flags such as organizations they had never heard of. Getting personalization right showed up in other categories as well. For example, many targets thought the technology communication medium (texting) was not appropriate for the type of message they received. Others recognized context inaccuracies, such as a colleague’s name that is made up. And while some people were more convinced by messages that had senders that were familiar, over two-thirds were turned off by inclusion of senders that were suspicious and not someone the person would know. All of these suggest the need for further research not only on what makes phishing messages persuasive, but how that interacts with personalization in spear phishing. We believe future studies can build upon the categories and insights related to spear phishing messages identified in this paper.
@@
This study illustrated the limitations of humans in trying to identify AI-generated spear phishing messages. Our quantitative results show that targets guessed accurately only 52% of the time, where 50% would be expected from randomly guessing. Qualitative results suggest that many individuals had no idea how to even approach this task, lacking mental models to know what criteria to use to differentiate them. This finding aligns with previous studies, such as those comparing AI-generated and human-written poetry, which demonstrate that people often struggle to differentiate between AI and human authorship [24]. Our results highlight a need for improved literacy and education for individuals against AI-driven phishing threats. One approach is to use LLMs to provide recommendations on how to deal with suspicious messages [22]. Another is to help users realize that they may expect to get more suspicious personalized messages in the future, given the lowering cost of creating spear phishing messages using AI. In the end, differentiating between AI-generated and human-generated messages is not as important as avoiding clicking on suspicious links.
Despite their lack of success, targets explained criteria they used to when guessing which messages were AI generated. Their lack of success and confidence in their guesses suggests that their criteria were more of superficial, heuristic cues when assessing the source of messages than hard and fast rules. Our results were consistent with previous research about human heuristics for AI-generated language [25]. Still, it is useful to better understand the perceived limitations that people put on AI. They described a variety of criteria including specific linguistic features (the presence of emojis, how the URLs look, word choice), as well as stylistic choices (length of message, content structure, grammar, and spelling). They commonly noted that AI-generated messages frequently exhibit peculiar word choices, overly formal or inconsistent styles, and unnatural content structures. Moreover, AI-generated texts were often perceived as either too perfect or containing subtle but noticeable grammatical and spelling errors. Also, messages that lacked specific personal details or felt generic were more likely to be identified as AI-generated, since people assumed that AI could not create highly personal content. This is particularly problematic, since this study demonstrates that, in fact, AI is likely better at creating highly personalized phishing messages than humans.

### v2

@@ 5.2.7 Plausible Rewards
Realistic rewards significantly influenced the perceived credibility of phishing messages, as noted by 28% of the participants (7 out of 25). One participant said they would be more likely to click on a link, “because the reward is connected to something I put so much time and effort into" (T31), while another participant stressed the persuasiveness of plausible or realistic offers, mentioning, “It’s like taking you to a link to look at an offer, but it’s not a crazy insane offer" (T16). Some targets expressed enthusiasm for realistic opportunities, stating, “It’s something I’ll be really excited about, but also something that I think is realistic" (T34). Conversely, 24% of participants (6 out of 25) identified messages with rewards that seemed “too good to be true" as less convincing. One participant mentioned it’s “probably not likely” to be the “developer of the month four times a year” (T37). One of the targets questioned the credibility of messages offering free access to goods and services, stating, “It sounds too good to be true" (T34). Another target warned against messages promoting anything for free, noting, “Anything that’s free is a little too good to be true. Okay, so that’s when I would be very careful" (T27). This collective sentiment emphasized the convincing nature of realistic rewards in phishing messages while avoiding extravagant or implausible claims that trigger suspicion.
@@ 5.2.8 Sender Familiarity
Participants described how the perceived credibility of phishing messages is tied to the sender’s identity, as noted by 7 out of 25 targets (28%). One participant said the message seemed more convincing because the sender “introduced herself as [their] neighbor" (T30). Another explained the importance of a personal introduction, stating, “The sender started with ’My name is John’" (T41), which helped the target to trust the message. Several participants expressed the importance of connecting with the senders of the messages. One target was convinced by the “actual company branding” that mimicked promotional texts they had previously received (T18). Another explained, “This could be someone from my ward [church congregation], wanting me to like check out some plants or something" (T31). Participants explained that when the message sender appears to be a legitimate, familiar source, messages seemed more credible.
Shifting towards participants’ suspicions, a substantial 68% (17 out of 25) highlighted the pivotal role of the sender in raising skepticism about received messages. One participant explained that messages appear suspicious when the introduction is not consistent with other messages from that same sender, saying, “James is my boss… it’d be weird that he was introducing himself that way… that doesn’t sound like my boss" (T31). Several participants described the need to understand how the sender unexpectedly obtained their phone number and the purpose for which the sender sent a message. One target noted, “So like, how did you get my number in the first place?" (T27), while another questioned, “Like why would Lowe’s have my phone number?" (T20). Participants described how recognition of the sender’s phone number is also an important component of message believability. One participant explained, “If it’s from a number I don’t recognize at all. Then yeah, probably the first thing I would ask is who it is. And then however, if the person is like, if their name is like the same as one of my friends, then I probably would open it" (T34). Participants described that messages lacking personal connections from the sender raised suspicions, with one stating, “I don’t feel like it’s anyone who’s personally connected with me; it just seems like spam, you know?" (T31). Being familiar with the message sender was described as a strong indicator in whether subjects believed the messages to be legitimate. It is important to note that none of the study text messages sent to participants indicated a source phone number. Rather, our focus was on the content of the messages.
@@ 5.3 Human/AI Source Identification
This subsection addresses RQ3, which asks how effective targets are at identifying AI-generated spear phishing SMS messages. The statistical analysis presented earlier showed that targets could not effectively identify AI-generated messages. The qualitative results presented here reinforce the difficulty of this task for targets. Although the targets described some of their reasoning behind their guesses on which messages are AI-generated (see following section), 12 out of 25 of the participants stated in some way that they were uncertain about their decisions, often relying on intuition rather than any specific criteria. One target remarked, “sometimes it’s a gut feeling maybe more than like a specific thing you’re looking for?" (T07), while another said it’s “just a feeling" (T02). Others attributed their lack of criteria to advancements in AI, saying that they,“don’t really, really know my criteria, because I know that AI is getting so good" (T04). Several targets openly admitted their lack of expertise, saying they “have no idea" (T15), and “actually I don’t know, I don’t know" (T23). Some were unsure if AI could personalize content or include icons in text (T02), while others conveyed concerns of AI’s competency, hoping that “the ones that are worse would hopefully be the AI ones" (T34). Overall, these responses show the difficulty the targets faced in distinguishing between AI and human-generated content. In general, they did not seem to have mental models that provided meaningful guidance in identifying AI-generated messages.
@@ 5.4 AI Identification Criteria
This subsection addresses RQ4 by identifying the criteria people use to try and identify AI-generated messages. Table 4 provides a summary of the percentage of targets who mentioned these criteria (or lack thereof) that influenced their determination of an AI-generated message.
Themes % Mentioned
No idea 48%
Style 40%
Emojis 20%
Personalization 32%
Word Choice 24%
Structure 24%
Grammar 24%
Message Length 16%
URL 8%
Table 4: Themes Influencing Perception of Messages as Human or AI-Generated. Although targets shared differing criteria for identifying the source of the message, a significant number do not have an idea.
5.4.1 Style
The message style was mentioned by 40% (10/25) of the targets. Many targets noted that AI-generated messages tend to be, “pretty formal" (T21) or “overly informal or overly formal" (T09) while others said messages often appear “extremely generic sounding," as if they were created from a prompt with specific parameters (T18). Additionally, the presence of “glaring flaws" such as typos and awkward phrasings were attributed to being human-made mistakes. Such common mistakes contrasted with the “more robotic" nature of AI messages (T08). Some targets observed that AI-generated messages might seem “too perfect," such as refraining from casual texting language that humans often use, such as “RN" for “right now" (T34). One participant noted that the excessive use of exclamation points is what made them think the message is AI generated: “I mean, honestly, what’s driving me insane about all of these is exclamation points. Yeah. I’m like, why are there so many exclamation points all over?" (T07). In contrast, messages that gave a sense of urgency were more likely to be perceived as human-written, as one target explained: “Then I didn’t put this one because they were actually pressuring me to do it immediately. So I don’t think AI can do it" (T02). Overall, targets indicated that a message’s style played a role in their AI vs human generated message decisions, with AI messages being attributed to using a style that is overly formal, generic, or laden with exclamation points; and with human messages appearing more casual, urgent, and imperfect.
@@ 5.4.2 Emojis
5.4.3 Personalization
The message personalization was mentioned by 32% (8/25) of the targets. Five participants mentioned that a lack of personalization made them suspect the message was AI-generated. For instance, one target noted, “AI were the ones that were a little bit less personal" (T34), while another expressed doubt, saying, “There is no connection with me as a reader" (T41). Some targets believed the absence of personal pronouns or a personal introduction with the recipient’s name indicated AI authorship. One target observed, “All of these messages, but two, include my name" (T37), leading them to suspect those two were AI-generated. Conversely, four participants cited personalized messages as evidence of human authorship, providing similar reasons as those who suspected AI. This group of participants felt messages tailored to the recipient sounded human, with one target commenting, “it just sounds very personal" (T21). Additionally, recipients noted that human-sounding messages often began with an introduction of the sender, such as “this is Sarah from the [my university] student instructor program" (T41). Overall, personalized messages were deemed more human-like, while those lacking personalization were associated with AI.
5.4.4 Word Choice
The message word choice was mentioned by 24% (6/25) of the targets. Four of the targets claimed that the word choice sounded like AI, particularly the use of buzzwords and marketing words. One target explained that messages sound AI-generated when they use buzzwords that are associated with being “attention grabbers” because “AI would utilize those a lot” (T16). Another target described a message as “too specific” when it included words that humans “wouldn’t necessarily use” when describing instructional design (T18). A few of the targets agreed that some of the messages had sentences that a human would have worded differently (T09). Conversely, two targets gave evidence of human-sounding messages and both mentioned slang words. One target described “There’s so many slang terms on it, that it seems really human to me” (T34), while another defended the messages they thought to be human-created and explained, “oftentimes, it was just because I felt like the language seemed a little more slang-like" (T06). Messages with specific, marketing-like words were perceived as AI-written while messages with casual, slang words that sounded more conversational were perceived as human-written.
5.4.5 Content/Structure
The structure of the message content was mentioned by 24% (6/25) of the targets. Participants pointed out various aspects of structure that hinted at AI involvement, including repetitive phrases and/or awkward sentence construction. For instance, one participant thought a message was AI generated because ““it looks like there’s a template. So there’s like a flow that you know, there’s a pattern you see it as maybe started by a human” (T41). Another noted instances where messages seemed generic or mass-produced, akin to “the email that is sent to everyone in the school" (T41). One also pointed out inconsistencies within messages, such as repetitive phrases or ““two different things, in the same message” (T04). One target also noted that AI-generated messages tended to be more wordy, containing “lots of filler phrases" compared to human-authored messages (T37).
5.4.6 Grammar/Spelling
5.4.7 Length of Message
The length of message was mentioned by 16% (4/25) of the targets. Participants assumed messages made by AI would be longer than human made messages. One mentioned that they would have “AI do the longer ones and I would write the shorter ones myself,” (T21) assuming those who created these messages would do the same. These four targets were correct in their hunch that AI messages would be longer. Across the messages, AI-generated messages have on average 41% higher character count per message, with 337.8 characters for AI, and 237.9 characters for human-authored messages.
@@ 6 Discussion
A growing body of literature has found evidence that AI-generated content from recent LLMs can outperform human-created content in various domains .[8, 29, 30, 31] For example, Zhang and Gosline compared content generated by GPT-4 with similar content created by professional content creators in the advertising field, finding that “Content generated by generative AI and augmented AI is perceived as of higher quality than that produced by human experts and augmented human experts" .[29] Nisbett and Spaiser explored the use of GPT-3 in creating moral statements supporting climate action, concluding that “GPT-3-generated statements are generally more convincing than human-generated statements" .[31] Similarly, in the domain of political speech, Palmer and Spirling found that “LLMs are capable of producing human-style arguments for different positions on subjects as varied as abortion, guns, immigration, and organ donation" and they could “out-perform human authors, though it varies by topic" .[30]
@@
Fewer studies have examined the ability of LLMs to create malicious content tailored to individuals, such as we examine in this paper. Heiding et al. examined phishing messages created by humans (using the V-triad approach), GPT-4, both human+GPT-4 combined, and a control message from an existing phishing dataset .[32] They found that the human V-triad and the combined human V-triad and GPT-4 approaches led to the highest click-through rate in a field experiment, followed by GPT-4 and then the control group.[32] Although they did not personalize these messages to each individual, they did personalize them to a specific university context. They also provide a cost analysis that demonstrates how inexpensive the creation of spear phishing messages can be. Hazell also points out that LLM-generated spear phishing messages can be “realistic" and “cost-effective," but does not provide systematic evidence comparing humans to AI, such as presented in this paper .[8]
@@
Our study contributes to this body of literature by comparing the effectiveness of AI-created and human-created spear phishing messages tailored to an individual target. We found that AI generally outperformed humans in creating spear phishing messages; however, the difference was not statistically significant, in part due to the relatively low sample size. As presented in Section 5.1.1, there is an 80% probability that GPT-4 is at least as good at or better than humans at creating these highly personalized spear phishing messages. While further research is justified in this area, we believe our findings provide a clear indication of the capabilities of LLMs for generating plausible spear phishing messages. It is important to note that these results are based on a very simple prompt, which could likely be improved. Furthermore, LLMs continue to improve in their quality, suggesting that AI will likely perform better at this task in the future. However, despite the fact that our human spear phishing authors had been trained on techniques to create spear phishing, it is possible that they could have been trained better. For example, they were not trained using the V-triad approach used so successfully by Heiding et al. .[32]
Participants in this study represented a range of job roles, including entry-level employees, mid-level managers, and executives across industries such as technology, healthcare, finance, and education. Despite this diversity, no significant differences were observed in the ability to differentiate AI-generated from human-generated messages, suggesting that the challenge of identifying these messages is consistent across demographic groups. However, qualitative responses indicated some role-specific variations in message evaluation. For instance, individuals in managerial positions found workplace-related messages more persuasive, while technical staff were more likely to scrutinize URLs and sender details. This reinforces the need for occupation-specific training to address diverse phishing vulnerabilities.
@@ 6.2 Characteristics of Convincing Spear Phishing Messages
Our statistical analysis found that targets considered job-related spear phishing messages twice as persuasive than messages related to hobbies or social media posts (38% intended click rate compared to 19% hobbies and 17% social media). This was true for both human and AI-generated messages. Spear phishing in the workplace carries significant risk, targeting both employees [33] and executives .[34] This suggests the need to be particularly vigilant for work-related spear phishing attacks. [35]
@@
The qualitative findings from this study point to several phishing message features that make them more convincing than others. [9, 12] Our findings aligned with prior research, revealing that messages matching the receiver’s expectations in terms of the sender, context, and relevance of a message were more convincing; those with poor grammar and spelling were less believable; and messages relating to the recipient’s life were more persuasive .[28, 36]
@@
We found that most of the features were either wholly persuasive or wholly unconvincing to all the targets. For example, personal relevance was completely correlated to the messages being persuasive while inaccuracies were completely correlated to the message being unconvincing. We also found that in some cases, some features may dissuade one person from clicking on a phishing message, while the same feature could influence another person to click. These findings were recently reported .[32] This suggests that some features may work better for certain individuals than others - a level of personalization that we have not yet examined, but seems possible to implement in future AI-generated spear phishing attacks that have more information about targets’ preferences. This personalization would be consistent with prior research. For example, one study found that people with different personality traits may be more susceptible to different types of attacks .[37]
@@
A key focus of this paper was on personalization of messages since these were spear phishing messages tailored to individuals. Personalizing messages can lead to more persuasive messages [38, 10, 6], but can also raise red flags when the personalization is even slightly off. In this study, over two-thirds of participants were likely to believe messages that related to topics of personal interest and relevance, showing the power of personalized phishing messages when done well .[39, 40] However, two-thirds of participants also shared examples of messages that were not believed because they included content that was not personally relevant or included red flags, such as referencing organizations they had never heard of. Getting personalization right showed up in other categories as well. For example, many targets thought the technology communication medium (texting) was not appropriate for the type of message they received. Others recognized context inaccuracies, such as a colleague’s name that is made up. And while some people were more convinced by messages that had familiar senders, over two-thirds were dissuaded by the inclusion of senders that were suspicious and not someone the person would know. All of these findings suggest the need for further research not only on what makes phishing messages persuasive, but how persuasion techniques relate to personalization in spear phishing .[39] We believe future studies can build upon the categories and insights related to spear phishing messages identified in this paper.
@@
This study illustrated the limitations of humans in trying to identify AI-generated spear phishing messages. Our quantitative results show that targets guessed accurately only 52% of the time, where 50% would be expected from randomly guessing. Qualitative results suggest that many individuals had no idea how to even approach this task, as they struggled to know what criteria to use to differentiate them. This finding aligns with previous studies, such as those comparing AI-generated and human-written poetry, which demonstrate that people often struggle to differentiate between AI and human authorship .[41] The inability to reliably differentiate AI- from human-generated messages is significant because it suggests that traditional heuristics for identifying phishing (e.g., grammar or spelling errors) may no longer apply. As AI-generated content becomes more refined, training individuals to detect subtle markers of malicious intent—rather than relying on assumed human traits—will be critical.
Despite their lack of success, targets explained criteria they used when guessing which messages were AI generated. Their lack of confidence when assessing the source of messages suggests that their assessment criteria were based more on superficial or heuristic cues rather than hard and fast rules. Our results were consistent with previous research about human heuristics for AI-generated language .[42] In this study, participants noted that the presence of first-person pronouns, use of contractions, and references to family topics were a sign of a human-authored message. However, these cues were not reliable and interestingly, the AI-generated texts that were optimized for perceived humanity were often rated as “more human” than actual human-written ones. Our results showed a similar pattern; many targets noted that AI has a writing style that is "overly formal", or "less personal" even though it is not a consistent or reliable marker for AI authorship.
Still, it is useful to better understand the perceived limitations that people put on AI. Participants described a variety of criteria including specific linguistic features (the presence of emojis, how the URLs look, word choice), as well as stylistic choices (length of message, content structure, grammar, and spelling). They commonly noted that AI-generated messages frequently exhibit peculiar word choices, overly formal or inconsistent styles, and unnatural content structures. Moreover, AI-generated texts were often perceived as either too perfect or containing subtle but noticeable grammatical and spelling errors. Also, messages that lacked specific personal details or felt generic were more likely to be identified as AI-generated, since people assumed that AI could not create highly personal content. These perceptions are particularly problematic, since this study demonstrates that, in fact, AI is likely better at creating highly personalized phishing messages than humans.
These insights highlight the dual importance of distinguishing AI-generated messages: (1) to identify their unique features for detection and (2) to improve awareness about their potential misuse. In the long term, recognizing the origin of messages could inform defenses (human or automated) against these emerging threats.